Testing: Ho: p-0.88 Hi:p +0.88 Your sample consists of 61 subjects, with 53 successes. Calculate the test statistic, rounded to 2 decimal places 2-
Answer #1

p^ = 53/61

z p1-p

   =\frac{53/61-0.88}{\sqrt{\frac{0.88*(1-0.88)}{61}}}

   = -0.27 (ans)